OpenStudy (anonymous):

ok this one im not sure of find the vertex, the line of symetry and the maximum or minimum value ofx and graph the function f(x)=-(x+9)^2-7

OpenStudy (anonymous):

Now to find Max an Min ...Just Differentiate them

OpenStudy (anonymous):

Maximum is (-9, -7), which is also the vertex. Line of symmetry is x = -9

OpenStudy (anonymous):

\[\frac{d f(x)}{dx} = -2(x +9) \] \[\frac{d^2 f(x)}{dx^2} = -2\]

OpenStudy (anonymous):

Then Put f'(x) = 0 -2(x +9) = 0 which gives you x = -9

OpenStudy (anonymous):

Now as f"(x) is minimum x =-9 turns to be maximum

OpenStudy (dumbcow):

i dont think he is in a calculus class guys. ok the quadratic equation is already in vertex form \[y = a(x-h)^{2} + k\] vertex = (h,k) when a is negative you have a downward parabola or a maximum

OpenStudy (anonymous):

can anyone graph this so i can see what it looks like?

OpenStudy (anonymous):

Now, The Function will form a Parabola and x =-9 should be line of symmetry

OpenStudy (anonymous):

It's just a downward opening parabola with the vertex at (-9, -7)

OpenStudy (dumbcow):

go to graphcalc.com

OpenStudy (anonymous):

ok so line of symetry is -9 and the maximum/minimum value is -7

OpenStudy (anonymous):

maximum is -9